The humidity level, X, in Tropicana Field in Tampa at the time of a night game ranges from 40% to 100 %. (.40 to 1.00) The rak) What is the probability that X is more than 0.1 above its expected value?
Answer #1

- giren CDF of arandom variablex is Ex(x)=10 ; nco-40 501-0.4)(2-31).airs 0.4 enclo 3 nalo a) The pdf of a random variable xi flan= 5(0.4-1) +5(2-2) 3 -- 5(2-2) + 5(2-0.4) 3 correct choice is b) The probability that x>015 is given by P(x>0.5] = 1- PEc P(x20.5] = 1-p (x20.5] =l- 0.75 = 0.25 from b) since x is a continuous random variable. p[x20.5) = P(x>0.5] Ed) P[x>0.6] =e> P[x>0.8] = 1-P[*5018] Ela 560-8-0.4)(2-08) = 15X0.4 x 1.2 = = 1-0.8 0.2 ATP[x20.8] =0.2 /